OpenStudy (anonymous):

Ralph wants to find the solution to a system of equations using y = x - 2 and y = x2 + 6x - 38. Joey says that Ralph can solve x2 + 5x - 36 = 0 to find the x-coordinates of the solutions to his system. Explain and demonstrate why Joey is correct.

OpenStudy (whpalmer4):

We have two equations: \[y = x-2\]\[y=x^2+6x-38\] Left side of both equations are identical. What does that imply about the right side of both equations?

OpenStudy (anonymous):

They could be set equal to each other?

OpenStudy (whpalmer4):

Yes! And what do you get if you do that, and then move everything to the same side of the equals sign?

OpenStudy (anonymous):

x-2=x^2+6x-38 +2 +2 x=x^2+6x-36 -x -x 0=x^2+5x-36 .. is that right?

OpenStudy (whpalmer4):

Yes.

OpenStudy (anonymous):

Okay thank youuu!